Help please
Correct Part G Enter the expression V2i +j+k where i. j and k are vectors. where i, j and k are vectors vec Select the vector and then click on the vedc button vec V2i+j+k (1.414,1,1) Submit Previous Answers Reguest Answer Provide Feedback O Type here to search
Answer #1

Given Vector is

V = \sqrt {2} i + j + k

\sqrt {2}i + j + k = \sqrt {2} \vec{i} + \vec{j} + \vec{k}

You need to use 'vec' button to write i, j, and k into vector form and use \sqrt to use the sign for square root.

Let me know if you have any query.